OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 3, 15, 75, 375 ... A. an=3(5)n-1 B. an=5(3)n-1 C. an=12(3)n-1 D. an=5(12)n-1

OpenStudy (***[isuru]***):

the general formula for a geometric sequence is .. \[A_{n} = ar^{n-1}\] where a = the first term r = the common ration the first term of the sequence is 3.. to find the common ration u have to divide 2 successive terms r = 15/3 = 5 so... the equation is \[A_{n} = 3. 5^{n-1}\] not it :) ?

OpenStudy (***[isuru]***):

*srry... i mean got .. not "not " XD

OpenStudy (anonymous):

yea, i got it. thanks .